CommercialServices.com Corporation provides business-to-business services on the Internet. Data concerning the most recent year appear below:

Sales $ 2,280,000
Net operating income $ 250,800
Average operating assets $ 760,000

The following questions are to be considered independently.

Garrison 16e Rechecks 2019-01-10

Required:

1. Compute the company's return on investment (ROI).

2. The entrepreneur who founded the company is convinced that sales will increase next year by 60% and that net operating income will increase by 100%, with no increase in average operating assets. What would be the company’s ROI?

3. The Chief Financial Officer of the company believes a more realistic scenario would be a $1,050,000 increase in sales, requiring a $572,000 increase in average operating assets, with a resulting $165,450 increase in net operating income. What would be the company’s ROI in this scenario?
Answer #1

1. ROI = net operating income / average operating assets

= 250800/760000 = 33%

2. ROI = (250800*2)/760000 = 66%

3. ROI = (250800+165450)/(760000+572000)

= 31.25%